# USTC-ICTS/PCFT-25-60

### v1

@@ 1 Introduction
The experimental discovery of the Higgs boson at the Large Hadron Collider (LHC) [1], which completes the full particle content predicted by the standard model (SM), has opened up a new era of precision measurements in particle physics. Of all the existing elementary particles, the top quark, the Higgs boson and the intermediate weak gauge bosons($W$ and $Z$) are the most massive ones, with masses around the electroweak scale $\sim{\cal O}(100)$ GeV. Thus the in-depth investigations, both experimentally and theoretically, of the properties of these heavy elementary particles will play an important role in precision SM tests and in looking for new physics beyond the SM.
Although the $W$ and $Z$ bosons were discovered more than forty years ago, only exclusive leptonic decays of theirs were observed, as shown by Particle Data Group [2], and no exclusive hadronic decays have been well established so far. It is of interest to study exclusive gauge boson decays into hadronic final states, which may present us with a new playground for precision electroweak and quantum chromodynamics (QCD) physics [3, 4, 5, 6].
Several years ago, the CMS Collaboration reported the first observation of rare $Z$-boson three-body decay, $Z\to\Psi\ell^{+}\ell^{-}$ ($\ell=e$ or $\mu$) [7], where $\Psi$ represents direct $J/\Psi$ and $J/\Psi$ from $\Psi(2S)\to J/\Psi X$ decays. After subtraction of the contribution from $\Psi(2S)$, the branching fraction ratio
@@ 2 Decay amplitudes
It is obvious that both ${\cal M}_{1}^{\gamma}$ and ${\cal M}_{2}$ are ${\cal O}(e^{2}g_{Z})$. Actually, if we replace the virtual photon in Figs. 1 and 2 by a virtual $Z$-boson, the corresponding amplitudes are ${\cal O}(g_{Z}^{3})$, which are of the same order in electroweak coupling constants as ${\cal M}_{1}^{\gamma}$ and ${\cal M}_{2}$. It will be shown below that, these ${\cal O}(g_{Z}^{3})$ amplitudes do not bring about significant contributions. However, in order to perform a complete calculation, we include both of them in the present study. It is easy to see that the amplitude contributed by $Z\to\ell^{+}\ell^{-}Z^{*}$, followed by $Z^{*}\to V$, can be incorporated into ${\cal M}_{1}^{\gamma}$, and their sum can be written in a compact form, which, denoted as ${\cal M}_{1}$, gives
@@
together with the total decay width of $Z$-boson $\Gamma_{Z}$, we can define the branching ratios as
@@
On the other hand, the analysis of the differential distributions with respect to $s$ or $\cos\theta$ will be also significant, which can be carried out by integrating over $\cos\theta$ or $s$ in eq. (14). Particularly, as will be shown below that, in the SM, the angular distribution is symmetric under $\cos\theta\leftrightarrow-\cos\theta$, consequently, no forward-backward asymmetries for final leptons can be expected. This implies that the forward-backward asymmetry, which is defined as
@@
could be an interesting observable to probe new physics beyond the SM.
@@ 3.1 In the SM
Obviously, except ${\cal B}_{1}$, others are negligible. Likewise, one can find that only ${\cal B}_{2}$ will get a little smaller and others remain almost unchanged for the dimuon final state. This is consistent with the statement in Refs. [14, 15] that, contributions from Fig. 2, relative to the ones from Fig. 1, are suppressed by the factor $m_{V}^{2}/m_{Z}^{2}$, which, for $J/\Psi$, is below the percent level or even smaller (In calculating ${\cal M}_{1}$, the contribution through $Z\to\ell^{+}\ell^{-}Z^{*}$, followed by $Z^{*}\to V$, has been included. Actually, this part can be neglected for the charmonium final states). Here we should emphasis that, to calculate the amplitude of Fig. 2, one has to introduce the non-perturbative quantity $\psi_{V}(0)$, which may not be determined very precisely; on the other hand, the non-perturbative effects due to strong interactions from Fig. 1, are confined to the matrix element (6) up to the decay constant $f_{V}$. As pointed out in Ref. [20], this approach has the advantage that it automatically takes into account higher order corrections to the electromagnetic current, which are common to both the electromagnetic decay and production of a vector meson. Since $f_{J/\Psi}$ can be fixed from the measured $J/\Psi\to e^{+}e^{-}$ width, a clean prediction for $Z\to J/\Psi\ell^{+}\ell^{-}$ with $\ell=e,~\mu$ can be expected. We thus obtain
@@
where the form of $\tilde{I}(s,\cos\theta)$ will be shown in eq. (3.2) in which $m_{J/\Psi}^{2}$ should be replaced by $m_{V}^{2}$. Therefore, the ratio of ${\cal B}(Z\to J/\Psi\ell^{+}\ell^{-})$ and ${\cal B}(Z\to\Psi(2S)\ell^{+}\ell^{-})$ can be determined up to the ratio of the experimental width of the charmonium leptonic decays, which reads
@@
Using the current experimental data, we have $R=3.24\pm 0.08$. Thus, the central value of ${\cal R}_{J/\Psi\ell^{+}\ell^{-}}$ given by the CMS Collaboration [7], shown in in eq. (1), in which $R=3.5$ was used, will be slightly shifted as ${\cal R}_{J/\Psi\ell^{+}\ell^{-}}=0.65$.
@@
for $Z\to\Upsilon(1S)e^{+}e^{-}$; while for the muon mode, only ${\cal B}_{2}$ will be changed as $4.29\times 10^{-10}$. One can easily find that ${\cal B}_{1}$ is still the most dominant, however, the other contributions from Fig. 2 via $Z\to V\gamma^{*}\to V\ell^{+}\ell^{-}$ can reach the percent level, and will be non-negligible any more. Meanwhile, the contributions via the virtual $Z$ transition remain insignificant. Totally, we have
@@
In the present work, in order to evaluate the diagrams in Fig. 2, we perform the computation in the framework of the nonrelativistic color-singlet model. One may use the nonrelativistic QCD (NRQCD) factorization method [22] to perform a systematical calculation in powers of $\alpha_{s}$ ($\alpha_{s}=g_{s}^{2}/4\pi$ and $g_{s}$ is the strong coupling constant) and $v$ ($v$ is the heavy-quark velocity in the quarkonium rest frame). Actually, our present computation based on the color-singlet model are equivalent to the one based on the NRQCD approach at the leading order. Therefore, to estimate the theoretical errors of our results, one can assume that the uncalculated QCD corrections in $\alpha_{s}$ are of order $\alpha_{s}(m_{V})$ and that the uncalculated corrections in $v$ are of order $v^{2}$. On the other hand, the calculation of ${\cal M}_{1}$ (then to $\Gamma_{1}$), which hadronizes the electromagnetic currents based on eq. (6), already contains the QCD and relativistic corrections to all orders. Thus the estimation of theoretical uncertainties from uncalculated corrections of order $\alpha_{s}$ and $v^{2}$ will be applied to $\Gamma_{i}$’s with $i\neq 1$. In the following, one will take
@@
Besides the branching ratios, one can further explore the differential distributions of $Z\to V\ell^{+}\ell^{-}$ decays. The invariant mass distributions for the lepton pair or the energy distributions for the vector mesons have been studied in Refs. [13, 14] in which only Fig. 1 was considered. For the charmonium case, since the contribution generated from Fig. 2 is can be neglected, these differential distributions should of course remain almost unchanged.
@@ 3.2 Beyond the SM
It is known that anomalous neutral gauge couplings (NTGC)[23, 24], which are absent in the SM, could induce CP-violating effects in $Z$-boson decays [25, 26]. The general form of the CP-violating NTGC involving the effective $Z(p)\gamma(q)G^{*}(k)$ vertex with $G=Z,~\gamma$ denoting the virtual gauge boson, has been given in Refs. [23, 24, 27], which reads
@@
Very recently, it has been claimed by the authors of Refs. [28] that the above conventional CP-violating NTGC is incompatible with the spontaneous breaking of the full electroweak gauge group, and a different structure for the second dimension-8 term of eq. (37) has been obtained in their work. In the present paper, we only consider the first dimension-6 $h_{1}^{G}$ term to carry out our analysis of the CP-violating effects in $Z\to V\ell^{+}\ell^{-}$ decays. Therefore, the new CP-violating amplitude can be straightforwardly written as
@@
${\cal A}(s)=\int^{1}_{-1}d\cos\theta~\beta_{\ell}~\lambda^{1/2}(m_{Z}^{2},m_{J/\Psi}^{2},s)~\tilde{I}(s,\cos\theta)$ (42)
with
$\tilde{I}(s,\cos\theta)=\left[\frac{2p\cdot k_{1}p\cdot k_{2}-m_{Z}^{2}k_{1}\cdot k_{2}}{(m_{J/\Psi}^{2}+2k_{1}\cdot q)^{2}}+\frac{2p\cdot k_{1}p\cdot k_{2}-m_{Z}^{2}k_{1}\cdot k_{2}}{(m_{J/\Psi}^{2}+2k_{2}\cdot q)^{2}}\right$.
$\left.+\frac{2k_{1}\cdot k_{2}(m_{Z}^{2}+m_{J/\Psi}^{2})}{(m_{J/\Psi}^{2}+2k_{1}\cdot q)(m_{J/\Psi}^{2}+2k_{2}\cdot q)}\right].$ (43)
Here we do not consider the contribution from the $h_{1}^{Z}$ term to ${\cal A}(s)$ since it is ${\cal O}[(h_{1}^{Z})^{2}]$ and negligible for the present range of $h_{1}^{Z}$ in (39). We have plotted $A_{\rm FB}(s)/h_{1}^{Z}$ with respected to $\sqrt{s}$, the dilepton invariant mass, in Fig. 6. Using the current bound on $h_{1}^{Z}$, the forward-backward asymmetry in $Z\to J/\Psi\ell^{+}\ell^{-}$, given by the anomalous $ZZ^{*}\gamma$ vertex, could be up to $10^{-3}$.
@@
which means the integrated asymmetry could reach the ${\cal O}(10^{-3})$ level. However, as we know, these decays are dominated by the electromagnetic fragmentation transition, the most events, as shown in Fig. 3, will intend to be in the large $s$ region. Thus, it is still not easy for experimental observations.

### v2

@@ 1 Introduction
The experimental discovery of the Higgs boson at the Large Hadron Collider (LHC) [1], which completes the full particle content predicted by the standard model (SM), has opened up a new era of precision measurements in particle physics. Of all the existing elementary particles, the top quark, the Higgs boson, and the intermediate weak gauge bosons ($W$ and $Z$) are the most massive ones, with masses around the electroweak scale $\sim{\cal O}(100)$ GeV. Thus the in-depth investigations, both experimentally and theoretically, of the properties of these heavy elementary particles will play an important role in precision SM tests and in looking for new physics beyond the SM.
Although the $W$ and $Z$ bosons were discovered more than forty years ago, only exclusive leptonic decays of theirs were observed, as shown by Particle Data Group [2], and no exclusive hadronic decays have been well established so far. It is of interest to study the exclusive gauge boson decaying into hadronic final states, which may present us with a new playground for precision electroweak and quantum chromodynamics (QCD) physics [3, 4, 5, 6].
Several years ago, the CMS Collaboration reported the first observation of a rare $Z$-boson three-body decay, $Z\to\Psi\ell^{+}\ell^{-}$ ($\ell=e$ or $\mu$) [7], where $\Psi$ represents direct $J/\Psi$ and $J/\Psi$ from $\Psi(2S)\to J/\Psi X$ decays. After subtraction of the contribution from $\Psi(2S)$, the branching fraction ratio
@@ 2 Decay amplitudes
It is obvious that both ${\cal M}_{1}^{\gamma}$ and ${\cal M}_{2}$ are ${\cal O}(e^{2}g_{Z})$. Actually, if we replace the virtual photon in Figs. 1 and 2 by a virtual $Z$-boson, the corresponding amplitudes are ${\cal O}(g_{Z}^{3})$, which are of the same order in electroweak coupling constants as ${\cal M}_{1}^{\gamma}$ and ${\cal M}_{2}$. It will be shown below that, these ${\cal O}(g_{Z}^{3})$ amplitudes do not bring about significant contributions. However, in order to perform a complete calculation, we include both of them in the present study. It is easy to see that the amplitude contributed by $Z\to\ell^{+}\ell^{-}Z^{*}$, followed by $Z^{*}\to V$, can be incorporated into ${\cal M}_{1}^{\gamma}$, and their sum can be written in a compact form, which, denoted as ${\cal M}_{1}$, gives  [^1] Now ${\cal M}_{1}$ looks like ${\cal O}(e^{2}g_{Z})$. However, the $Z^{*}\to V$ amplitude can be restored as ${\cal O}(g_{Z}^{3})$ by using the relation $e^{2}/\sin^{2}2\theta_{W}\sim g_{Z}^{2}$ in the second term of $\tilde{g}_{V}^{\ell}$ and $\tilde{g}_{A}^{\ell}$ in eq. (11).
@@
together with the total decay width of $Z$-boson $\Gamma_{Z}=2.4955$ GeV [2], we can define the branching ratios as
@@
On the other hand, the analysis of the differential distributions with respect to $s$ or $\cos\theta$ will be also significant, which can be carried out by integrating over $\cos\theta$ or $s$ in eq. (14). Particularly, as shown in the Appendix, the differential rate $d^{2}\Gamma_{1}/ds~d\cos\theta$ is the even function of $\cos\theta$, and the same conclusion can be reached for other differential rates $d^{2}\Gamma_{i}/ds~d\cos\theta$ with $i\neq 1$. Thus, the angular distribution is symmetric under $\cos\theta\leftrightarrow-\cos\theta$, consequently, no forward-backward asymmetries for final leptons can be expected in the SM. This implies that the forward-backward asymmetry, which is defined as
@@
could be an interesting observable to probe new physics beyond the SM. In the next section, we will analyze this quantity by introducing the anomalous neutral gauge coupling terms. It will be shown below, due to the new interaction, the linear term of $\cos\theta$ could appear in the differential distributions, which may lead to nonzero $A_{\rm FB}$ in $Z\to V\ell^{+}\ell^{-}$ decays.
@@ 3.1 In the SM
Obviously, except ${\cal B}_{1}$, others are negligible since their sum is below the percent level relative to ${\cal B}_{1}$ (For instance, the ratio ${\cal B}_{2}/{\cal B}_{1}$ is about $0.2\%$). Likewise, one can find that only ${\cal B}_{2}$ will get a little smaller and others remain almost unchanged for the dimuon final state. This is consistent with the statement in Refs. [14, 15] that, contributions from Fig. 2, relative to the ones from Fig. 1, are suppressed by the factor $m_{V}^{2}/m_{Z}^{2}$, which, for $J/\Psi$, is below the percent level or even smaller (In calculating ${\cal M}_{1}$, the contribution through $Z\to\ell^{+}\ell^{-}Z^{*}$, followed by $Z^{*}\to V$, has been included. Actually, this part can be neglected for the charmonium final states). Here we wish to emphasize that, to calculate the amplitude of Fig. 2, one has to introduce the non-perturbative quantity $\psi_{V}(0)$, which may not be determined very precisely; on the other hand, the non-perturbative effects due to strong interactions from Fig. 1, are confined to the matrix element (6) up to the decay constant $f_{V}$. As pointed out in Ref. [20], this approach has the advantage that it automatically takes into account higher order corrections to the electromagnetic current, which are common to both the electromagnetic decay and production of a vector meson. Since $f_{J/\Psi}$ can be fixed from the measured $J/\Psi\to e^{+}e^{-}$ width, a clean prediction for $Z\to J/\Psi\ell^{+}\ell^{-}$ with $\ell=e,~\mu$ can be expected. We thus obtain
@@
where the form of $\tilde{I}(s,\cos\theta)$ is explicitly shown in eq. (Appendix: Some analytic expressions of the differential decay rate) of the Appendix. Therefore, the ratio of ${\cal B}(Z\to J/\Psi\ell^{+}\ell^{-})$ to ${\cal B}(Z\to\Psi(2S)\ell^{+}\ell^{-})$ can be determined up to the ratio of the experimental width of the charmonium leptonic decays, which reads
@@
Using the current experimental data, we have $R=3.24\pm 0.08$. Thus, the central value of ${\cal R}_{J/\Psi\ell^{+}\ell^{-}}$ given by the CMS Collaboration [7], shown in eq. (1), in which $R=3.5$ was used, will be slightly shifted as ${\cal R}_{J/\Psi\ell^{+}\ell^{-}}=0.65$.
@@
for $Z\to\Upsilon(1S)e^{+}e^{-}$; while for the muon mode, only ${\cal B}_{2}$ will be changed as $4.29\times 10^{-10}$. One can easily find that ${\cal B}_{1}$ is still the most dominant, however, the other contributions from Fig. 2 via $Z\to V\gamma^{*}\to V\ell^{+}\ell^{-}$ can reach the percent level, and will be non-negligible any more (For instance, the ratio ${\cal B}_{2}/{\cal B}_{1}$ is about $5\%$). Meanwhile, the contributions via the virtual $Z$ transition remain insignificant. Totally, we have
@@
In the present work, in order to evaluate the diagrams in Fig. 2, we perform the computation in the framework of the nonrelativistic color-singlet model. Here we would like to give some remarks on the uncertainties of our theoretical predictions from this approach.
First, the non-perturbative quantity $\psi_{V}(0)$, which is taken as an input, cannot be determined precisely. Its uncertainty has been included in our numerical calculation, whose effects are negligible for the bottomonium final states. As to the charmonium case, the contribution from Fig. 2 can be neglected.
Second, one may go beyond the present approach by using the nonrelativistic QCD (NRQCD) factorization method [22] to perform a systematical calculation in powers of $\alpha_{s}$ ($\alpha_{s}=g_{s}^{2}/4\pi$ and $g_{s}$ is the strong coupling constant) and $v$ ($v$ is the heavy-quark velocity in the quarkonium rest frame). Actually, our present computation based on the color-singlet model are equivalent to the one based on the NRQCD approach at the leading order. Therefore, to estimate the theoretical errors of our results, one can assume that the uncalculated QCD corrections in $\alpha_{s}$ are of order $\alpha_{s}(m_{V})$ and that the uncalculated corrections in $v$ are of order $v^{2}$. On the other hand, the calculation of ${\cal M}_{1}$ (then to $\Gamma_{1}$), which hadronizes the electromagnetic currents based on eq. (6), already contains the QCD and relativistic corrections to all orders. Thus the estimation of theoretical uncertainties from uncalculated corrections of order $\alpha_{s}$ and $v^{2}$ will be applied to $\Gamma_{i}$’s with $i\neq 1$. In the following, one will take
@@
Besides the branching ratios, one can further explore the differential distributions of $Z\to V\ell^{+}\ell^{-}$ decays. The invariant mass distributions for the lepton pair or the energy distributions for the vector mesons have been studied in Refs. [13, 14] in which only Fig. 1 was considered. For the charmonium case, since the contribution generated from Fig. 2 can be neglected, these differential distributions should of course remain almost unchanged.
@@ 3.2 Beyond the SM
It is known that anomalous neutral triple gauge couplings (NTGC)[23, 24], which are absent in the SM, could induce CP-violating effects in $Z$-boson decays [25, 26]. The general form of the CP-violating NTGC involving the effective $Z(p)\gamma(q)G^{*}(k)$ vertex with $G=Z,~\gamma$ denoting the virtual gauge boson, has been given in Refs. [23, 24, 27], which reads
@@
Very recently, it has been claimed by the authors of Refs. [28] that the above conventional CP-violating NTGC is incompatible with the spontaneous breaking of the full electroweak gauge group, and a different structure for the second dimension-8 term of eq. (37) has been obtained in their work. In the present paper, we only consider the first dimension-6 $h_{1}^{G}$ term to carry out our analysis of the CP-violating effects in $Z\to V\ell^{+}\ell^{-}$ decays. [^2] Generally speaking, the second dimension-8 term would be more suppressed by the high scale of new physics. However, its contribution might be important in some new scenarios. In the present work, for simplicity, we just choose the dimension-6 term to illustrate our numerical analysis. Therefore, the new CP-violating amplitude can be straightforwardly written as
@@
${\cal A}(s)=\int^{1}_{-1}d\cos\theta~\beta_{\ell}~\lambda^{1/2}(m_{Z}^{2},m_{J/\Psi}^{2},s)~\tilde{I}(s,\cos\theta).$ (42)
The expression of $\tilde{I}(s,\cos\theta)$ can be found in eq. (Appendix: Some analytic expressions of the differential decay rate) of the Appendix in which $m_{V}^{2}$ should be taken as $m_{J/\Psi}^{2}$. Here we do not consider the contribution from the $h_{1}^{Z}$ term to ${\cal A}(s)$ since it is ${\cal O}[(h_{1}^{Z})^{2}]$ and negligible for the present range of $h_{1}^{Z}$ in (39). Obviously, using eqs. (A5) and (A6), the term $q\cdot(k_{2}-k_{1})$ in (3.2) can give rise to the linear term of $\cos\theta$. We have plotted $A_{\rm FB}(s)/h_{1}^{Z}$ with respect to $\sqrt{s}$, the dilepton invariant mass, in Fig. 6. Using the current bound on $h_{1}^{Z}$, the forward-backward asymmetry in $Z\to J/\Psi\ell^{+}\ell^{-}$, given by the anomalous $ZZ^{*}\gamma$ vertex, could be up to $10^{-3}$.
@@
which means the integrated asymmetry could reach the ${\cal O}(10^{-3})$ level. [^3] The decay rate of this process is dominated by the electromagnetic fragmentation transition, which is in the large $s$ region. Thus, in the above cut region, the denominator of eq. (43) will be substantially decreased, which can enhance the integrated asymmetry. However, as we know, these decays are dominated by the electromagnetic fragmentation transition, the most events, as shown in Fig. 3, will intend to be in the large $s$ region. Thus, it is still not easy for experimental observations.
@@ Acknowledgments
Appendix: Some analytic expressions of the differential decay rate
The full expression of $Z\to V\ell^{+}\ell^{-}$ decay at the tree-level in the SM can be derived using eq. (14) together with eqs. (9), (10), (12) and (13). Here we take $\Gamma_{1}$} as an example to show our derivation explicitly.
$\frac{d^{2}\Gamma_{1}}{ds~d\cos\theta}$ $=$ $\frac{1}{512\pi^{3}m_{Z}^{3}}\beta_{\ell}~\lambda^{1/2}(m_{Z}^{2},m_{V}^{2},s)~\frac{1}{3}\sum_{\rm spins}|{\cal M}_{1}|^{2}$ (A1)
$=$ $\frac{g_{Z}^{2}\Gamma(V\to e^{+}e^{-})}{32\pi^{2}m_{Z}^{2}m_{V}}[(\tilde{g}_{V}^{\ell})^{2}+(\tilde{g}_{A}^{\ell})^{2}]\beta_{\ell}\lambda^{1/2}(m_{Z}^{2},m_{V}^{2},s)\tilde{I}(s,\cos\theta)$
with
$\tilde{I}(s,\cos\theta)=\left[\frac{2p\cdot k_{1}p\cdot k_{2}-m_{Z}^{2}k_{1}\cdot k_{2}}{(m_{V}^{2}+2k_{1}\cdot q)^{2}}+\frac{2p\cdot k_{1}p\cdot k_{2}-m_{Z}^{2}k_{1}\cdot k_{2}}{(m_{V}^{2}+2k_{2}\cdot q)^{2}}\right$.
$\left.+\frac{2k_{1}\cdot k_{2}(m_{Z}^{2}+m_{V}^{2})}{(m_{V}^{2}+2k_{1}\cdot q)(m_{V}^{2}+2k_{2}\cdot q)}\right].$ (A2)
On the other hand, one can easily derive
$p\cdot k_{1}=\frac{m_{Z}^{2}-m_{V}^{2}+s}{4}-\frac{1}{4}\beta_{\ell}\lambda^{1/2}(m_{Z}^{2},m_{V}^{2},s)\cos\theta,$ (A3)
$p\cdot k_{2}=\frac{m_{Z}^{2}-m_{V}^{2}+s}{4}+\frac{1}{4}\beta_{\ell}\lambda^{1/2}(m_{Z}^{2},m_{V}^{2},s)\cos\theta,$ (A4)
$m_{V}^{2}+2k_{1}\cdot q=\frac{m_{Z}^{2}+m_{V}^{2}-s}{2}-\frac{1}{2}\beta_{\ell}\lambda^{1/2}(m_{Z}^{2},m_{V}^{2},s)\cos\theta,$ (A5)
$m_{V}^{2}+2k_{2}\cdot q=\frac{m_{Z}^{2}+m_{V}^{2}-s}{2}+\frac{1}{2}\beta_{\ell}\lambda^{1/2}(m_{Z}^{2},m_{V}^{2},s)\cos\theta.$ (A6)
By further neglecting the lepton mass, we have $k_{1}\cdot k_{2}=s/2$. This shows that the function $\tilde{I}$ and the differential rate (A1) can be in terms of the kinematical variables $s$ and $\cos\theta$ completely. It is straightforward to see that this differential rate is the even function of $\cos\theta$. One can similarly deal with the rest parts in eq. (14).